# A Methodology for Incompleteness-Tolerant and Modular Gradual Semantics for Argumentative Statement Graphs

### v6

@@ 4.2 Properties
Given a second SG $\mathcal{G}^{\prime}\!=\!\langle\mathcal{X}\!\cup\!\{\alpha_{1}\},\mathcal{A}^{\prime},\mathcal{S}^{\prime},\tau^{\prime}\rangle$, a GS $\sigma$ satisfies neutrality iff for any $\alpha_{2}\!\in\!\mathcal{X}$, where $\mathcal{A}^{\prime}(\!\alpha_{2}\!)\!\cup\!\mathcal{S}^{\prime}(\!\alpha_{2}\!)\!\!=\!\!\mathcal{A}(\!\alpha_{2}\!)\!\cup\!\mathcal{S}(\!\alpha_{2}\!)\!\cup\!\{\alpha_{1}\}$, $\sigma^{\mathcal{G}^{\prime}}(\!\alpha_{3}\!)\!=\!\sigma^{\mathcal{G}}(\!\alpha_{3}\!)$ $\forall\alpha_{3}\in\mathcal{A}(\!\alpha_{2}\!)\!\cup\!\mathcal{S}(\!\alpha_{2}\!)$, $\tau^{\prime}(\!\alpha_{2}\!)\!=\!\tau(\!\alpha_{2}\!)$ and $\sigma^{\mathcal{G}^{\prime}}(\!\alpha_{1}\!)\!=\!0$, $\sigma^{\mathcal{G}^{\prime}}(\!\alpha_{2}\!)\!=\!\sigma^{\mathcal{G}}(\!\alpha_{2}\!)$.
This property, as with stability, takes a notably different approach to the T-norm semantics given their focus on CSTs, namely that the strengths of statements are calculated recursively, but should be unaffected by bottom-strength statements.
@@
Given a second SG $\mathcal{G}^{\prime}=\langle\mathcal{X}\cup\{\alpha_{1}\},\mathcal{A}^{\prime},\mathcal{S}^{\prime},\tau^{\prime}\rangle$, a GS $\sigma$ satisfies attacked premise iff for any $\alpha_{2}\in\mathcal{X}$, where $\mathcal{A}^{\prime}(\alpha_{2})\!=\!\mathcal{A}(\alpha_{2})\cup\{\alpha_{1}\}$, $\mathcal{S}^{\prime}(\alpha_{2})\!=\!\mathcal{S}(\alpha_{2})$, $\sigma^{\mathcal{G}^{\prime}}(\alpha_{3})\!=\!\sigma^{\mathcal{G}}(\alpha_{3})$ $\forall\alpha_{3}\!\in\!\mathcal{A}(\alpha_{2})\cup\mathcal{S}(\alpha_{2})$ and $\tau^{\prime}(\alpha_{2})\!=\!\tau(\alpha_{2})$, $\sigma^{\mathcal{G}^{\prime}}(\alpha_{2})\!\leq\!\sigma^{\mathcal{G}}(\alpha_{2})$.
@@
Given a second SG $\mathcal{G}^{\prime}=\langle\mathcal{X}\cup\{\alpha_{1}\},\mathcal{A}^{\prime},\mathcal{S}^{\prime},\tau^{\prime}\rangle$, a GS $\sigma$ satisfies supported premise iff for any $\alpha_{2}\in\mathcal{X}$, where $\mathcal{A}^{\prime}(\alpha_{2})\!=\!\mathcal{A}(\alpha_{2})$, $\mathcal{S}^{\prime}(\alpha_{2})\!=\!\mathcal{S}(\alpha_{2})\!\cup\!\{\alpha_{1}\}$, $\sigma^{\mathcal{G}^{\prime}}(\alpha_{3})\!=\!\sigma^{\mathcal{G}}(\alpha_{3})$ $\forall\alpha_{3}\!\in\!\mathcal{A}(\alpha_{2})\cup\mathcal{S}(\alpha_{2})$ and $\tau^{\prime}(\alpha_{2})\!=\!\tau(\alpha_{2})$, $\sigma^{\mathcal{G}^{\prime}}(\alpha_{2})\!\geq\!\sigma^{\mathcal{G}}(\alpha_{2})$.
@@
Given a second SG $\mathcal{G}^{\prime}=\langle\mathcal{X}^{\prime},\mathcal{A}^{\prime},\mathcal{S}^{\prime},\tau^{\prime}\rangle$, a GS $\sigma$ satisfies weakened premise iff for any $\alpha_{1}\in\mathcal{X}$, where $\mathcal{A}^{\prime}(\alpha_{1})=\mathcal{A}(\alpha_{1})$, $\mathcal{S}^{\prime}(\alpha_{1})=\mathcal{S}(\alpha_{1})$, $\tau^{\prime}(\alpha_{1})=\tau(\alpha_{1})$, and $\sigma^{\mathcal{G}^{\prime}}(\alpha_{2})=\sigma^{\mathcal{G}}(\alpha_{2})$ $\forall\alpha_{2}\in\mathcal{A}(\alpha_{1})\cup\mathcal{S}(\alpha_{1})\setminus\{\alpha_{3}\}$, where $\alpha_{3}\in\mathcal{A}(\alpha_{1})$ and $\sigma^{\mathcal{G}^{\prime}}(\alpha_{3})>\sigma^{\mathcal{G}}(\alpha_{3})$, $\sigma^{\mathcal{G}^{\prime}}(\alpha_{1})\leq\sigma^{\mathcal{G}}(\alpha_{1})$.
@@
Given a second SG $\mathcal{G}^{\prime}=\langle\mathcal{X}^{\prime},\mathcal{A}^{\prime},\mathcal{S}^{\prime},\tau^{\prime}\rangle$, a GS $\sigma$ satisfies strengthened premise iff for any $\alpha_{1}\in\mathcal{X}$, where $\mathcal{A}^{\prime}(\alpha_{1})=\mathcal{A}(\alpha_{1})$, $\mathcal{S}^{\prime}(\alpha_{1})\!=\!\mathcal{S}(\alpha_{1})$, $\tau^{\prime}(\alpha_{1})=\tau(\alpha_{1})$, and $\sigma^{\mathcal{G}^{\prime}}(\alpha_{2})=\sigma^{\mathcal{G}}(\alpha_{2})$ $\forall\alpha_{2}\in\mathcal{A}(\alpha_{1})\cup\mathcal{S}(\alpha_{1})\setminus\{\alpha_{3}\}$, where $\alpha_{3}\in\mathcal{S}(\alpha_{1})$ and $\sigma^{\mathcal{G}^{\prime}}(\alpha_{3})>\sigma^{\mathcal{G}}(\alpha_{3})$, $\sigma^{\mathcal{G}^{\prime}}(\alpha_{1})\geq\sigma^{\mathcal{G}}(\alpha_{1})$.
These properties assess added statements’ effects, conditioning on their strengths rather than weights and direct attackers and supporters rather than CSTs, thus accommodating statements of all three levels of completeness.
@@
For partially-complete or incomplete statements especially, this property provides an alternative reason for requiring rejection, i.e., there is maximal evidence to the contrary of one of its premises since they are, critically, part of a conjunction, as opposed to provability, where this is the case unless every premise in the statement is supported. Note that bottom-strength premise is also desirable for complete statements.
@@
In our running example, let $\alpha_{10}=\langle\neg a,\neg e\rangle$ (“CO2 emissions are not rising, thus extreme weather events are not increasing”) with $\tau(\alpha_{5})\!=\!\tau(\alpha_{10})\!=\!0.5$. Then, $\sigma^{\mathcal{G}}_{\wedge_{D}}(\alpha_{10})\!=\!1-\sigma^{\mathcal{G}}_{\wedge_{D}}(\alpha_{5})\!=\!1-0.55\!=\!0.45$, and $\sigma^{\mathcal{G}}_{T_{p}}(\alpha_{5})\!=\!0$ and $\sigma^{\mathcal{G}}_{T_{p}}(\alpha_{10})\!=\!0.45$. Despite $\alpha_{5}$ and $\alpha_{10}$ having premises that negate one another and equal base weights, their strengths are not complementary (do not sum to 1) under T-norm-p semantics. In contrast, the DC semantics show a desirable duality in attackers’ and supporters’ effects.
@@ 4.3 Theoretical Analysis
2: Rewriting $\checkmark$ $\checkmark$ $\times$ $\times$ -- --
3: Provability $\checkmark$ $\checkmark$ $\times$ $\times$ -- --
4: Weak Provability $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ -- --
@@
11: Bottom-Strength Premise $\checkmark$ $\checkmark$ $\checkmark$ $\times$ -- --
12: Top-Strength Premises $\times$ $\times$ $\checkmark$ $\times$ -- --
13: Mirroring $\times$ $\times$ $\checkmark$ $\checkmark$ -- --
Table 1: Structured ($\sigma_{T_{p}}$, $\sigma_{T_{m}}$, $\sigma_{\wedge_{D}}$ and $\sigma_{\wedge_{Q}}$) and abstract ($\sigma_{D}$ and $\sigma_{Q}$) GS’ satisfaction (denoted with $\checkmark$) or violation (denoted with $\times$) of the assessed properties, where incompatibility of a GS with a property is denoted with --.
@@
Both GS satisfy identical properties across our collection, however, they perform more poorly when we consider the novel properties. Even stability and neutrality, two basic properties in the realm of abstract GS, are violated due to the fact that they do not consider the weight as a starting point, nor the strength as a recursive function, respectively. Attacked and supported premises are both satisfied, while only weakened, and not strengthened, premise is satisfied due to the asymmetry in the way the attackers and supporters are evaluated, which leads to mirroring being violated also. Similarly, bottom-strength premises is satisfied, but top-strength premises is violated due to the (potentially non-maximal) weight of the statement playing a role in the calculation. Given top-strength premises’ fundamental role across statements of any completeness, this represents a significant weakness.
@@
This means that a user selecting a semantics must choose between its satisfation of one set of properties or another, so our approach represents a viable alternative to the existing work, particularly in cases with incomplete information. Examples 3 and 4 demonstrate what we believe is further justification for this. For example, rewriting is not satisfied by either semantics since more evidence, even if the statements constituting it are not complete, will strengthen an statement, which seems to be reasonable particularly for our settings with partially-complete or incomplete statements. Moreover, provability is violated by design, since we wanted GS that do not necessarily reject a statement when its supporting evidence is missing, only doing so when the statement’s weight is also minimal, hence the semantics’ satisfaction of weak provability instead. Both GS align with stability in treating the weight as a “starting point”, and from there they satisfy attacked and supported premises, as well as weakened and strengthened premises, due to the recursive way they handle attackers and supporters. Interestingly, the DC semantics, when supported by the DF-QuAD semantics, satisfies bottom-strength premise and top-strength premises, but it does not when supported by the QEM semantics. This is due to the saturation effect that is seen at the extremities of the strength scale in the DF-QuAD semantics, giving rise to this intuitive behaviour for this specific type of SG (with conjunctions as premises). Finally, mirroring is satisfied by both GS.
@@ 5 Conclusions and Future Work
Note that while this paper establishes the foundations of our novel approach using a toy example for pedagogical reasons, we see significant potential for real-world applications, particularly in the analysis of debates mapped onto SGs through argument mining (?), possibly enhanced by large language models (??). It is well known that mining structured arguments is a complex problem, not least due to the presence of enthymemes causing incompleteness (??). A methodology leveraging GS that can handle incomplete information could alleviate some of the burden that structured argumentation’s requirements for completeness place on these methods.
Our methodology represents an early step forward in GS for structured argumentation, and opens various potentially fruitful directions for future work. First, we would like to broaden our analysis, both within our restricted form of SGs, exploring various other instantiations of our DC semantics, e.g., using other abstract GS such as the exponent-based restricted semantics (?), as well as extending to different forms of SGs with more complex logical premises. Specifically, our ability to deal with partially-complete or incomplete statements may be related to the possibility of deriving assumptions in non-flat ABA (?). We would also like to consider whether other properties for abstract GS are desirable, e.g., open-mindedness (?) (the satisfaction of which gives a clear advantage of $\sigma_{Q}$ over $\sigma_{D}$), attainability (?), whose suitability in our restricted SGs was not obvious due to the conjunctive premise. It would also be interesting to formalise the interplay between properties, e.g., between provability and stability, which are incompatible without restrictions. Further, it would be interesting to see whether our analysis could be applicable to probabilistic argumentation (?????). Finally, we plan to explore how our methodology translates to real-world applications, e.g., in multi-agent model reconciliation (?) given GS’ capability for modelling human or machine reasoning (?).

### v7

@@ 4.2 Properties
Given two SGs $\mathcal{G}$ and $\mathcal{G}^{\prime}\!=\!\langle\mathcal{X}\!\cup\!\{\alpha_{1}\},\mathcal{A}^{\prime},\mathcal{S}^{\prime},\tau^{\prime}\rangle$, a GS $\sigma$ satisfies neutrality iff for any $\alpha_{2}\!\in\!\mathcal{X}$, where $\mathcal{A}^{\prime}(\!\alpha_{2}\!)\!\cup\!\mathcal{S}^{\prime}(\!\alpha_{2}\!)\!\!=\!\!\mathcal{A}(\!\alpha_{2}\!)\!\cup\!\mathcal{S}(\!\alpha_{2}\!)\!\cup\!\{\alpha_{1}\}$, $\sigma^{\mathcal{G}^{\prime}}(\!\alpha_{3}\!)\!=\!\sigma^{\mathcal{G}}(\!\alpha_{3}\!)$ $\forall\alpha_{3}\in\mathcal{A}(\!\alpha_{2}\!)\!\cup\!\mathcal{S}(\!\alpha_{2}\!)$, $\tau^{\prime}(\!\alpha_{2}\!)\!=\!\tau(\!\alpha_{2}\!)$ and $\sigma^{\mathcal{G}^{\prime}}(\!\alpha_{1}\!)\!=\!0$, $\sigma^{\mathcal{G}^{\prime}}(\!\alpha_{2}\!)\!=\!\sigma^{\mathcal{G}}(\!\alpha_{2}\!)$.
This property, as with stability, takes a notably different approach to the T-norm semantics given their focus on CSTs, namely that the strengths of statements are calculated recursively, but should be unaffected by bottom-strength statements, which must be unsubstantiated, as illustrated below.
@@
Given two SGs $\mathcal{G}$ and $\mathcal{G}^{\prime}=\langle\mathcal{X}\cup\{\alpha_{1}\},\mathcal{A}^{\prime},\mathcal{S}^{\prime},\tau^{\prime}\rangle$, a GS $\sigma$ satisfies attacked premise iff for any $\alpha_{2}\in\mathcal{X}$, where $\mathcal{A}^{\prime}(\alpha_{2})\!=\!\mathcal{A}(\alpha_{2})\cup\{\alpha_{1}\}$, $\mathcal{S}^{\prime}(\alpha_{2})\!=\!\mathcal{S}(\alpha_{2})$, $\sigma^{\mathcal{G}^{\prime}}(\alpha_{3})\!=\!\sigma^{\mathcal{G}}(\alpha_{3})$ $\forall\alpha_{3}\!\in\!\mathcal{A}(\alpha_{2})\cup\mathcal{S}(\alpha_{2})$ and $\tau^{\prime}(\alpha_{2})\!=\!\tau(\alpha_{2})$, $\sigma^{\mathcal{G}^{\prime}}(\alpha_{2})\!\leq\!\sigma^{\mathcal{G}}(\alpha_{2})$.
@@
Given two SGs $\mathcal{G}$ and $\mathcal{G}^{\prime}=\langle\mathcal{X}\cup\{\alpha_{1}\},\mathcal{A}^{\prime},\mathcal{S}^{\prime},\tau^{\prime}\rangle$, a GS $\sigma$ satisfies supported premise iff for any $\alpha_{2}\in\mathcal{X}$, where $\mathcal{A}^{\prime}(\alpha_{2})\!=\!\mathcal{A}(\alpha_{2})$, $\mathcal{S}^{\prime}(\alpha_{2})\!=\!\mathcal{S}(\alpha_{2})\!\cup\!\{\alpha_{1}\}$, $\sigma^{\mathcal{G}^{\prime}}(\alpha_{3})\!=\!\sigma^{\mathcal{G}}(\alpha_{3})$ $\forall\alpha_{3}\!\in\!\mathcal{A}(\alpha_{2})\cup\mathcal{S}(\alpha_{2})$ and $\tau^{\prime}(\alpha_{2})\!=\!\tau(\alpha_{2})$, $\sigma^{\mathcal{G}^{\prime}}(\alpha_{2})\!\geq\!\sigma^{\mathcal{G}}(\alpha_{2})$.
@@
Given two SGs $\mathcal{G}$ and $\mathcal{G}^{\prime}=\langle\mathcal{X}^{\prime},\mathcal{A}^{\prime},\mathcal{S}^{\prime},\tau^{\prime}\rangle$, a GS $\sigma$ satisfies weakened premise iff for any $\alpha_{1}\in\mathcal{X}$, where $\mathcal{A}^{\prime}(\alpha_{1})=\mathcal{A}(\alpha_{1})$, $\mathcal{S}^{\prime}(\alpha_{1})=\mathcal{S}(\alpha_{1})$, $\tau^{\prime}(\alpha_{1})=\tau(\alpha_{1})$, and $\sigma^{\mathcal{G}^{\prime}}(\alpha_{2})=\sigma^{\mathcal{G}}(\alpha_{2})$ $\forall\alpha_{2}\in\mathcal{A}(\alpha_{1})\cup\mathcal{S}(\alpha_{1})\setminus\{\alpha_{3}\}$, where $\alpha_{3}\in\mathcal{A}(\alpha_{1})$ and $\sigma^{\mathcal{G}^{\prime}}(\alpha_{3})>\sigma^{\mathcal{G}}(\alpha_{3})$, $\sigma^{\mathcal{G}^{\prime}}(\alpha_{1})\leq\sigma^{\mathcal{G}}(\alpha_{1})$.
@@
Given two SGs $\mathcal{G}$ and $\mathcal{G}^{\prime}=\langle\mathcal{X}^{\prime},\mathcal{A}^{\prime},\mathcal{S}^{\prime},\tau^{\prime}\rangle$, a GS $\sigma$ satisfies strengthened premise iff for any $\alpha_{1}\in\mathcal{X}$, where $\mathcal{A}^{\prime}(\alpha_{1})=\mathcal{A}(\alpha_{1})$, $\mathcal{S}^{\prime}(\alpha_{1})\!=\!\mathcal{S}(\alpha_{1})$, $\tau^{\prime}(\alpha_{1})=\tau(\alpha_{1})$, and $\sigma^{\mathcal{G}^{\prime}}(\alpha_{2})=\sigma^{\mathcal{G}}(\alpha_{2})$ $\forall\alpha_{2}\in\mathcal{A}(\alpha_{1})\cup\mathcal{S}(\alpha_{1})\setminus\{\alpha_{3}\}$, where $\alpha_{3}\in\mathcal{S}(\alpha_{1})$ and $\sigma^{\mathcal{G}^{\prime}}(\alpha_{3})>\sigma^{\mathcal{G}}(\alpha_{3})$, $\sigma^{\mathcal{G}^{\prime}}(\alpha_{1})\geq\sigma^{\mathcal{G}}(\alpha_{1})$.
These properties assess added statements’ effects, conditioning on their strengths rather than weights, and direct attackers and supporters rather than CSTs, thus accommodating statements of all three levels of completeness.
@@
For partially-complete or incomplete statements especially, this property provides an alternative reason for requiring rejection, i.e., there is maximal evidence to the contrary of one of its premises since they are, critically, part of a conjunction, as opposed to provability, where this is the case unless every premise in the statement is supported. Note that bottom-strength premise is also desirable for complete statements. We illustrate this property in the following.
@@
In our running example, let $\alpha_{10}=\langle\neg a,\neg e\rangle$ (“CO2 emissions are not rising, thus extreme weather events are not increasing”) with $\tau(\alpha_{5})\!=\!\tau(\alpha_{10})\!=\!0.5$. Then, $\sigma^{\mathcal{G}}_{\wedge_{D}}(\alpha_{10})\!=\!1-\sigma^{\mathcal{G}}_{\wedge_{D}}(\alpha_{5})\!=\!1-0.55\!=\!0.45$, and $\sigma^{\mathcal{G}}_{T_{p}}(\alpha_{5})\!=\!0$ and $\sigma^{\mathcal{G}}_{T_{p}}(\alpha_{10})\!=\!0.45$. Despite $\alpha_{5}$ and $\alpha_{10}$ having premises that negate one another and equal base weights, their strengths are not complementary (do not sum to 1) under T-norm-p semantics. In contrast, the DC semantics shows a desirable duality in attackers’ and supporters’ effects.
@@ 4.3 Theoretical Analysis
2: Rewriting $\checkmark$ $\checkmark$ $\times$ $\times$ - -
3: Provability $\checkmark$ $\checkmark$ $\times$ $\times$ - -
4: Weak Provability $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ - -
@@
11: Bottom-Strength Premise $\checkmark$ $\checkmark$ $\checkmark$ $\times$ - -
12: Top-Strength Premises $\times$ $\times$ $\checkmark$ $\times$ - -
13: Mirroring $\times$ $\times$ $\checkmark$ $\checkmark$ - -
Table 1: Structured ($\sigma_{T_{p}}$, $\sigma_{T_{m}}$, $\sigma_{\wedge_{D}}$ and $\sigma_{\wedge_{Q}}$) and abstract ($\sigma_{D}$ and $\sigma_{Q}$) GS’ satisfaction (denoted with $\checkmark$) or violation (denoted with $\times$) of the assessed properties, where incompatibility of a GS with a property is denoted with -.
@@
Both GS satisfy identical properties across our collection, performing poorly when we consider the novel properties. Even stability and neutrality, two basic properties in the realm of abstract GS, are violated due to the fact that they do not consider the weight as a starting point, nor the strength as a recursive function, respectively. Attacked and supported premises are both satisfied, while only weakened, and not strengthened, premise is satisfied due to the asymmetry in the way the attackers and supporters are evaluated, which leads to mirroring being violated also. Similarly, bottom-strength premises is satisfied, but top-strength premises is violated due to the (potentially non-maximal) weight of the statement playing a role in the calculation. Given top-strength premises’ fundamental role across statements of any completeness, this represents a significant weakness.
@@
This means that a user selecting a semantics must choose between its satisfation of one set of properties or another, so our approach represents a viable alternative to the existing work, particularly in cases with incomplete information. Examples 3 and 4 demonstrate what we believe is further justification for this. For example, rewriting is not satisfied by either semantics since more evidence, even if the statements constituting it are not complete, will strengthen a statement, which seems to be reasonable particularly for our settings with partially-complete or incomplete statements. Moreover, provability is violated by design, since we wanted GS that do not necessarily reject a statement when its support is missing, only when the statement’s weight is also minimal, hence the GS’ satisfaction of weak provability instead. Both GS align with stability in treating the weight as a “starting point”, and from there they satisfy attacked and supported premises, as well as weakened and strengthened premises, due to the recursive way they handle attackers and supporters. Interestingly, the DC semantics, when supported by the DF-QuAD semantics, satisfies bottom-strength premise and top-strength premises, but it does not when supported by the QEM semantics. This is due to the saturation effect that is seen at the extremities of the strength scale in the DF-QuAD semantics, giving rise to this intuitive behaviour for this specific type of SG (with conjunctions as premises). Finally, mirroring is satisfied by both GS.
@@ 5 Conclusions and Future Work
While this paper establishes the foundations of our novel approach using a toy example for pedagogical reasons, we see significant potential for real-world applications, e.g., in the analysis of debates mapped onto SGs through argument mining (?), possibly enhanced by large language models (??). It is well known that mining structured arguments is a complex problem, not least due to the presence of enthymemes causing incompleteness (??). A methodology leveraging GS that can handle incomplete information could alleviate some of the burden that structured argumentation’s requirements for completeness place on these methods.
Our methodology represents an early step forward in GS for structured argumentation, and opens various potentially fruitful directions for future work, in particular targeting more general settings of structured argumentation, e.g., in ABA or ASPIC+. First, we would like to broaden our analysis, both within our restricted form of SGs, exploring various other instantiations of our DC semantics, e.g., using other abstract GS such as the exponent-based restricted semantics (?), as well as extending to different forms of SGs with more complex logical premises. Specifically, our ability to deal with partially-complete or incomplete statements may be related to the possibility of deriving assumptions in non-flat ABA (?). We would also like to consider whether other properties for abstract GS are desirable, e.g., open-mindedness (?) (the satisfaction of which gives a clear advantage of $\sigma_{Q}$ over $\sigma_{D}$), attainability (?), whose suitability in our SGs was not obvious due to the conjunctive premise. It would also be interesting to formalise the interplay between properties, e.g., between provability and stability, which are incompatible without restrictions. Further, it would be interesting to investigate whether our analysis could be applicable to probabilistic argumentation (?????). Finally, we plan to explore how our methodology translates to real-world applications, e.g., in multi-agent model reconciliation (?) given GS’ capability for modelling human or machine reasoning (?).
Acknowledgements
Antonio Rago and Francesca Toni are partially funded by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 101020934) and by J.P. Morgan and the Royal Academy of Engineering under the Research Chairs and Senior Research Fellowships scheme (grant agreement no. RCSRF2021\11\45). Stylianos Loukas Vasileiou and William Yeoh are partially supported by the National Science Foundation (NSF) under award 2232055. Son Tran is partially supported by the NSF grants #1914635, #2151254, and ExpandAI grant #2025-67022-44266. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the sponsoring organizations, agencies, or governments.